STRUCTURE OF NAPHTHALENE DIMER INFERRED FROM SPECTRA OF NAPHTHALENE DIMER - ARGON VAN DER WAALS COMPLEXES

NEERAJ SHARMA, CARINE GILLIÉRON, B. KLAAS NAUTA and \underline{TIMOTHY W. SCHMIDT}, School of Chemistry, University of Sydney, NSW 2006, Australia.

We have obtained resonant 2-colour 2-photon ionization spectra of Np_{n}Ar_{m} (Np = naphthalene). While most spectra are broad and unresolved, those of Np_{2}Ar_{m} arouse particular attention as they appear to consist of two independent and shifted electronic spectra. We consider our experimental data in the light of proposed structures for the naphthalene dimer and offer an alternative explanation for the observed broadening of the spectrum of Np_{2} in terms of fragmentation of higher clusters.